OpenStudy (anonymous):

solve: 5a-3/2-3a-9/8=6

OpenStudy (jack1):

is this: \[\frac{ 5a-3 }{ 2-3a }-9/8\] or is it: \[\frac{ 5a-3 }{ 2 }-\frac{ 3a-9 }{ 8 }\]

OpenStudy (anonymous):

the question is quite confusing, i suggest you use parentheses :D

OpenStudy (anonymous):

Could you show us what you've done so far?

OpenStudy (dwelsher):

5/a-3/2-3a-9=6 2a=3/2+9/8+6 2a=12/8+9/8+6 2a=21/8+48/8 2a=69/8 16a=69 a=4.3125

OpenStudy (texaschic101):

5a - 3/2 - 3a - 9/8 = 6 (multiply by 8 to get rid of fractions) (8)5a - (8)3/2 - (8)3a - (8)9/8 = (8)6 40a - 12 - 24a - 9 = 48 16a - 21 = 48 16a = 48 + 21 16a = 69 a = 69/16 a = 4.3